OpenStudy (anonymous):

Solve by factoring: show all work 4x^2 + 8x − 5 = 0

OpenStudy (anonymous):

4x^2 +10x -2x -5 =0 2x (2x +5) -1(2x +5) =0 (2x +5) (2x -1) =0 Solve 2x +5 = 0 or 2x -1 =0 x = -5/2 or x = 1/2

OpenStudy (anonymous):

4x^2 + 8x − 5 = 0 4x^2 +10x-2x − 5 = 0 2x (2x +5) -1(2x +5) =0 (2x +5) (2x -1) =0 2x +5 = 0 or 2x -1 =0 x = -5/2 or x = 1/2

hero (hero):

That's the same thing the other person did, lol

OpenStudy (anonymous):

yes your right

hero (hero):

Nick, how do we know you didn't just copy and paste?

OpenStudy (anonymous):

why would i do that

OpenStudy (anonymous):

sorry you wasn't taking about me

OpenStudy (anonymous):

but your right it looks the same